#345 AT Decant - US Unit

From a US Unit perspective, this project started at the beginning of April with regular "Appleton Tower Decant Working Group" meetings. See 303 - AT/FH Decant Final Report for more details, including CO time figures. A decant schedule was prepared in advance and was agreed with Teaching Office and the Working Group.

As noted in the decant report above, there were a number of constraints to the move which meant we had a very tight schedule to work to. We weren't able to start the decant until 18th May with a firm deadline of 15th June to have vacated AT. We employed 2 student helpers for a total of 10 weeks. Without this effort, we would not have been able to meet the deadline and maintain an acceptable level of support in the Forum without using additional help from other Units which we wanted to avoid.

The deadline of 15th June for decanting from AT was met. Thereafter the vast majority of our effort was spent on setting up ISS and student labs in FH, the remainder being taken up by our own moves to Wilkie. The target date of 1st June for having the Drill Hall (1.B30) available to students was met as far as having machines in place. However there was a delay to opening the Drill Hall due to Health and Safety concerns and the unreliability of the air-conditioning. A big thank you must go to those involved in setting up the Drill Hall as they had to work in extremely unpleasant high temperatures. All other labs were installed by the end of July (with 2 labs on the third floor being upgraded to SL7).

Issues

There were no major show-stoppers but a few issues along the way:

  • Communication - principally with the project managers from Estates and Buildings - we were not invited to meetings without being pro-active on our part and even when we raised issues in advance no action was taken until the last minute
  • Excessive heat in Drill Hall and in the new 'support office'.
    • The temperature in the Drill Hall has been rectified. Although the temperature in the support office has improved, it still remains unpleasant.
  • ISIS broke 2 flip-desks in transit from AT to FH. These were subsequently replaced but the existing 'bins' didn't fit. This took a lot of chasing to be resolved.
  • It wasn't clear until very late on exactly how intrusive the work would be. We had hoped to be able to leave items in the switch cupboards but this turned out not to be feasible resulting in a last minute flurry of activity to move items to the secure store on level 6 AT.
  • 2 new desks delivered and installed by Estates and Buildings were left with no tops! This was eventually resolved after chasing.
  • We installed over 100 new 24" monitors in the Drill Hall and adjacent labs. Due to the design of the new monitors, we weren't able to loop the security cable through them so had to purchase new locks which attach to the back of the monitors.

Effort

  • Student helpers, 10 weeks
  • CSO, 12 weeks

-- AlisonDownie - 13 Aug 2015

Topic revision: r8 - 07 Oct 2015 - 09:19:52 - TimColles
 
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback
This Wiki uses Cookies